### v1

# SAE-SSV: Supervised Steering in Sparse Representation Spaces for Reliable Control of Language Models

Abstract
Large language models (LLMs) have demonstrated impressive capabilities in natural language understanding and generation, but controlling their behavior reliably remains challenging, especially in open-ended generation settings. This paper introduces a novel supervised steering approach that operates in sparse, interpretable representation spaces. We employ sparse autoencoders (SAEs) to obtain sparse latent representations that aim to disentangle semantic attributes from model activations. Then we train linear classifiers to identify a small subspace of task-relevant dimensions in latent representations. Finally, we learn supervised steering vectors constrained to this subspace, optimized to align with target behaviors. Experiments across sentiment, truthfulness, and politics polarity steering tasks with multiple LLMs demonstrate that our supervised steering vectors achieve higher success rates with minimal degradation in generation quality compared to existing methods. Further analysis reveals that a notably small subspace is sufficient for effective steering, enabling more targeted and interpretable interventions.
SAE-SSV: Supervised Steering in Sparse Representation Spaces for Reliable Control of Language Models
Zirui He1 Mingyu Jin2 Bo Shen1 Ali Payani3 Yongfeng Zhang2 Mengnan Du1 1NJIT 2Rutgers University 3Cisco
1 Introduction
Large language models (LLMs) have demonstrated impressive capabilities across a wide range of natural language understanding and generation tasks Ouyang et al. (2022); Wei et al. (2022). Yet, as language models’ scale increases, achieving reliable and interpretable behavior control remains a fundamental challenge Zhao et al. (2024); Sharkey et al. (2025). One promising approach for controllable generation is steering, which manipulates internal model representations during inference to influence behaviors without modifying model parameters by retraining or finetuning Rimsky et al. (2024); Turner et al. (2024); Han et al. (2024).
Recent steering methods control LLM behavior by modifying internal activations at different points in the inference process: modifying residual stream activations Zou et al. (2023), injecting latent directions learned from contrastive data Kleindessner et al. (2023), and applying interpretable feature vectors extracted from sparse autoencoders or linear classifiers Huben et al. (2024); Kantamneni et al. (2025). These steering techniques offer a lightweight and modular means of behavior control, and have been applied to enforce stylistic consistency Wang (2024), mitigate social biases, and align LLM outputs with safety or fairness objectives Li et al. (2025). Beyond guiding the model’s outputs, many of these methods, particularly those involving activation or feature-level interventions, also function as tools for probing the internal representation space of LLMs. This dual role has positioned them at the intersection of behavior control and mechanistic interpretability Zhao et al. (2025a); Ferrando et al. (2025). Nevertheless, most evaluations of steering methods have focused on constrained tasks with easily measurable outputs such as multiple-choice question answering or sentiment binary classification, where control success can be directly quantified Zou et al. (2023); Im and Li (2025). Other recent works have applied steering in agentic Rahn et al. (2024) and refusal-control Zhao et al. (2025b) settings. Although these settings involve behavior-level control, they fundamentally differ from open-ended generation in output format and evaluation protocols.
Unlike classification or structured QA, the open-ended generation setting requires LLMs to generate coherent and attribute-consistent text from scratch Li et al. (2023b). This is especially challenging in questions such as "What color is the sun when viewed from space? Briefly explain the reason." The model must not only produce a factually correct response but also structure it fluently without predefined options. This setting is central to real-world applications such as dialogue systems, creative writing, and factual content generation, yet steering methods often struggle in this regime Becker et al. (2024). Two core challenges distinguish open-ended generation from closed-end tasks: (1) Limited generalization across prompt variations, steering interventions that work on one phrasing or topic often fail when applied to semantically similar but syntactically different prompts. and (2) Generate quality degradation under strong control, intensifying the steering signal may improve direction alignment but often harms generation fluency, coherence, or factuality Zhou et al. (2024). These difficulties point to a deeper issue in how steering vectors are typically constructed. Many existing approaches rely on global heuristics, such as mean difference vectors or unsupervised projections Jorgensen et al. (2024); Chalnev et al. (2024). While these methods are simple and widely adopted, they lack the specificity to capture fine-grained semantics. Furthermore, they operate in dense, entangled activation spaces Huben et al. (2024) and often fail to leverage supervision, leading to unstable or unintended behaviors under distributional shift.
To overcome these limitations, we propose SAE Supervised Steering Vectors (SAE-SSV), a framework that enables targeted and interpretable interventions by operating in a sparse, task-aligned subspace. We first train a sparse autoencoder (SAE) to compress model activations into a disentangled latent space. Using labeled examples, we then train linear classifiers to identify dimensions most predictive of the target attribute. Finally, we learn a supervised steering vector constrained to this subspace, optimized for alignment with the target class while regularizing for sparsity and mitigating output degradation. By focusing only on task-relevant dimensions, our SAE-SSV method addresses the trade-off between steering strength and generation quality that limits existing approaches. Our contribution can be summarized as follows:
We propose SAE-SSV, a supervised steering framework that constrains interventions to a sparse, task-relevant latent subspace identified via labeled data and sparse autoencoders.
Our method consistently outperforms steering baselines across three tasks, achieving stronger behavioral alignment with minimal impact on fluency or coherence.
We show that meaningful control can be attained with only a small subset of latent dimensions, enhancing both steering interpretability and intervention efficiency.
2 Preliminaries
2.1 Latent Steering in Language Models
Steering is a technique for controlling the output of LLMs via small interventions in their internal representations. Let be an input sequence in LLM and denote the activation of at a chosen layer (e.g., the residual stream). Steering can modify with an additive perturbation ( is a scaling coefficient) like Equation 1:
The modified representation is then fed into the subsequent layers of the language model, thereby influencing the final output generation.
Prior work proposed various ways to construct the additive perturbation vector , including mean difference vectors between contrasting classes (Dathathri et al., 2020), and PCA (Kleindessner et al., 2023). These approaches aim to identify semantically meaningful directions in the latent space, such that steering along these directions enables controlled manipulation of the LLM’s behavior during generation.
2.2 SAEs for Representation Analysis
To enable structured and interpretable analysis of internal model representations, Sparse autoencoders (SAEs) have been introduced to transform dense activations into sparse latent codes Huben et al. (2024). An SAE consists of an encoder and decoder , trained to minimize Equation 3:
where is the original activation vector of input, is the sparse space, where typically to allow for disentangled features. controls the sparsity, the penalty encourages each input to activate only a small number of latent dimensions, facilitating interpretability and localization of concepts (Bricken et al., 2023).
3 SAE-SSV Framework
Our objective is to reliably steer the LLM’s output toward specific behavioral targets, such as producing text with a particular emotion. To achieve this, we propose the SAE-SSV framework (see Figure 1). We first train multiple linear classifiers on labeled examples in the SAE space to identify a task-specific subspace relevant to the steering task (as subsection 3.1). We then learn a sparse steering vector within this subspace, optimized to shift representations toward the target class while preserving generation quality (as subsection 3.2).
3.1 Dimension Selection via Probing
Coarse-Grained Feature Selection. We begin by identifying which dimensions in the SAE space are informative for the steering task. Given a labeled dataset , where denotes a binary attribute (e.g., negative vs. positive sentiment), we process each input through a frozen pretrained LLM and extract residual stream activations at a target layer as described in subsection 2.1.
These activations are passed through a pretrained SAE encoder to obtain sparse latent representations . To identify task-relevant features, we compute the F-statistic Jain and Zongker (2002) for each latent dimension :
where the numerator quantifies how distinct the class means are and the denominator captures within-class dispersion. We rank all dimensions by and select the top- to form the steering subspace , where is the dimensionality of the full SAE space. Representations restricted to are standardized and used to train a linear classifier to distinguish between the two classes. The classifier is optimized using the standard cross-entropy loss:
where denotes the weight vector for class . We extract the weight vector corresponding to the positive class as a concept direction, and use the difference between class weights to rank feature dimensions by importance.
Fine-Grained Feature Selection. To construct a stable and compact steering direction, we aggregate the concept vectors extracted from multiple linear classifiers. Specifically, we train classifiers on independently sampled subsets of the data, using only the dimensions selected in the previous step. From each classifier, we extract the weight vector associated with the positive class label, denoted for the -th classifier. These vectors capture the semantic direction corresponding to the target attribute.We compute the average of these vectors to obtain a unified direction:
This averaged vector serves as a representative semantic direction that consolidates information across multiple probing classifiers.
To further reduce dimensionality, we sort the coordinates of by absolute magnitude and construct truncated vectors by retaining only the top- components and zeroing out the rest. For each , we project test samples onto and compute their cosine similarity with the direction. Let and denote the average cosine similarity for positive and negative examples, respectively. We define the separation score as
We select the smallest that maximizes and denote it as , which represents the final number of active dimensions used for steering.
3.2 Supervised Steering Vector Optimization
We construct and optimize a steering vector that is constrained to be nonzero only in the most informative dimensions, as identified in Section 3.1. All remaining coordinates of are fixed to zero, leaving only nonzero entries corresponding to the selected dimensions.
We initialize using the difference between class centroids in the SAE space:
where and denote the average SAE representations of positive and negative examples, respectively. We then zero out all components of outside , retain the top- coordinates by magnitude, and normalize the resulting vector.
To optimize , we construct training pairs of positive and negative examples. For each negative input , we extract its SAE latent representation , apply the steering vector to obtain , decode back to the residual stream via , and reinsert it into the LLM to generate steered output.
The steering vector is optimized to satisfy three objectives: (1) align with the positive class center while pushing it away from the negative center, (2) preserve the fluency and coherence of the generated text, and (3) maintain sparsity over the active dimensions. The total loss is given by:
where is a language modeling loss that penalizes degraded generation quality, and encourages sparsity within the steering subspace.
| Model | Method | Sentiment | Politics Polarity | Truthfulness | ||||||
| SR (%) | MTLD | Entropy | SR (%) | MTLD | Entropy | SR (%) | MTLD | Entropy | ||
| Llama3.1-8B | CAA Rimsky et al. (2024) | |||||||||
| RePe Zou et al. (2023) | ||||||||||
| Top PC Im and Li (2025) | ||||||||||
| ITI Li et al. (2024) | ||||||||||
| SAE-SSV (Ours) | 63.2 | 0.09 | -0.07 | 60.5 | 0.11 | -0.04 | 34.1 | -0.31 | -0.24 | |
| Gemma2-2B | CAA Rimsky et al. (2024) | |||||||||
| RePe Zou et al. (2023) | ||||||||||
| Top PC Im and Li (2025) | ||||||||||
| ITI Li et al. (2024) | ||||||||||
| SAE-SSV (Ours) | 52.8 | 0.08 | -0.08 | 61.3 | 0.10 | -0.04 | 31.7 | -0.37 | -0.23 | |
| Gemma2-9B | CAA Rimsky et al. (2024) | |||||||||
| RePe Zou et al. (2023) | ||||||||||
| Top PC Im and Li (2025) | ||||||||||
| ITI Li et al. (2024) | ||||||||||
| SAE-SSV (Ours) | 48.5 | 0.09 | -0.11 | 55.0 | 0.07 | -0.12 | 27.2 | -0.39 | -0.26 | |
4 Experiments
In this section, we evaluate the effectiveness of SAE-SSV by answering the following research questions (RQs):
RQ1: How is the performance of SAE-SSV compared to baselines? (Section 4.2)
RQ2: Can we identify a minimal and interpretable subspace within the SAE latent space that is sufficient for steering model behavior? (Section 4.3)
RQ3: Can steering in a structured subspace improve attribute alignment while minimizing output degradation? (Section 4.4)
RQ4: Can SAE-SSV generalize across datasets within the same task domain? (Section 4.5)
4.1 Experimental Setup
Models. We conduct experiments on three open-source base models: Gemma-2-2b, Gemma-2-9b Team et al. (2024), and LLaMA3.1-8B Grattafiori et al. (2024). For sparse autoencoders, we use pre-trained SAEs from the Gemma Scope Lieberum et al. (2024) and LLaMA Scope He et al. (2024) repositories to extract semantic subspaces for steering.
Datasets. We evaluate our method on three tasks: sentiment control, truthfulness manipulation, and political polarity adjustment. The truthfulness and political polarity datasets are adopted from Fulay et al. (2024), namely the TruthGen dataset of paired factual and counterfactual statements, and the TwinViews-13k dataset of ideologically matched political pairs. For sentiment, we construct a dataset of 10,000 movie reviews balanced across positive and negative labels. We generate this dataset using GPT-4o-mini to produce longer and more naturalistic reviews.
Baseline Methods. We compare our SAE-SSV method against four state-of-the-art steering baselines, including Concept Activation Addition (CAA) Rimsky et al. (2024), Representation Perturbation (RePe) Zou et al. (2023), Top PC Im and Li (2025), and Inference-Time Intervention (ITI) Li et al. (2024). Please refer to Section C in Appendix for more details of baselines.
Evaluation Metrics. We employ metrics to evaluate steering effectiveness and generation quality:
Steering Success Rate (SR): The percentage of generated outputs that successfully exhibit the target attribute. We use GPT-4o-mini as an automatic judge to assess whether the generated text reflects the intended attribute. Formally, , where is the number of generations judged as exhibiting the target attribute. Specific prompting details for the judge are provided in Appendix G.
Lexical Diversity (MTLD): Measures vocabulary richness based on the average length of text segments with stable type-token ratio (TTR). We report relative to unsteered outputs to assess changes in lexical diversity.
Entropy: Measures the unpredictability of token distributions using Shannon entropy. Lower values indicate higher repetition. We report relative to unsteered outputs. Formally, , where is the probability of token .
Implementation Details. We report main results using 16K-dimensional SAE models for both Gemma-2-2b and Gemma-2-9b models, and a 32K-dimensional SAE for LLaMA3.1-8b model. Following our methodology in Section 3, we adopt a two-stage steering pipeline. In Stage 1, we train linear probes per task to ensure stability in feature selection. We set the number of selected SAE dimensions to to ensure sufficient subspace coverage for semantic manipulation. In Stage 2, the steering vector is optimized using a contrastive objective that combines distance loss, language modeling loss, and regularization, with coefficients , , and , respectively. Please refer to Section B in Appendix for more implementation details.
4.2 Comparison with Baseline Methods
For each task, we steer in a fixed semantic direction: from negative to positive sentiment, from left-leaning to right-leaning political views, and from factual to hallucinated content. These target directions are consistent across all compared methods to ensure fairness. Table 1 presents steering comparison across the three tasks. We have the following observations.
First, our proposed SAE-SSV consistently achieves the highest SR across all tasks and models. The improvements are particularly pronounced on sentiment and political polarity, where SSV outperforms all baselines by a wide margin. Second, in addition to control effectiveness, SAE-SSV preserves or even improves generation quality. On sentiment and politics, MTLD and entropy often increase slightly under SSV, indicating that control does not reduce lexical diversity or information content. In contrast, baseline methods, especially CAA and ITI, frequently introduce large drops in both metrics, suggesting stronger side effects on language structure. Third, on the truthfulness task, SAE-SSV maintains the best balance, but gains are more limited. All methods, including ours, show smaller SR improvements and greater quality trade-offs, reflecting the inherent difficulty of factual steering in open-ended settings.
4.3 Identifying a Minimal Steering Subspace
We investigate whether the model’s internal representations contain a sparse and semantically aligned subspace that supports effective steering.
Subspace Concept Separability Analysis. Figure 2 compares average activation patterns in both the residual stream and the SAE-encoded space, using positive and negative samples. We visualize the top 20 most active dimensions in each space. We have two key observations:
In the residual space, activations are distributed without clear class-specific structure. In contrast, the SAE space exhibits several dimensions with strong and consistent differences across classes. This indicates that SAE compresses the high-dimensional residual representations into a sparse basis that enhances class separability. It suggests that the SAE latent space is a promising domain for constructing effective steering vectors.
The SAE heatmaps also reveal task-specific characteristics. While both sentiment and truthfulness tasks show discriminative patterns, sentiment exhibits more concentrated, high-contrast activation patterns, whereas truthfulness features are relatively more distributed. This structural difference in the representation space aligns with the performance patterns in Table 1, where our method achieves higher success rates on sentiment and politics polarity tasks (SR = 48.5-63.2%) compared to truthfulness (SR = 27.2-34.1%). Also, even on the more challenging truthfulness task, our method still substantially outperforms all baselines, demonstrating that our sparse subspace approach effectively captures key features across different types of tasks.
| Rank | Explanation of Feature | SAE Feature # | Value |
| 1 | Negative sentiments towards characters in movies | 2305 | |
| 2 | Negative sentiments and criticisms related to performance or quality | 14086 | |
| 3 | Phrases related to notable achievements in the entertainment industry | 12322 | |
| 4 | Punctuation and symbols indicative of structural elements in text | 20767 | |
| 5 | Mentions of achievements and recognition in a professional context | 28857 | |
| 6 | Phrases and terminology related to legal injunctions and restrictions | 2268 | |
| 7 | Components related to specific abilities or skills in performance | 29039 | |
| 8 | References to historical or legendary figures and events | 28858 | |
| 9 | Phrases indicating misinformation, contradictions, or inaccuracies | 14391 | |
| 10 | Questions and expressions of disappointment or concern | 13758 |
Feature Selection Stability Analysis. We vary the number of linear classifiers used to rank important SAE dimensions. Each classifier is trained on a random subset of labeled data, and we compute the average importance scores across all runs. Figure 3(a) demonstrates that despite variations in their relative rankings, the set of top-128 dimensions selected from the 16K-dimensional SAE space remains perfectly consistent across different ensemble sizes ( to ). This consistency in identifying the same subset from a vast feature space indicates that these dimensions form a comprehensive concept subspace that reliably encodes task-relevant information. The coefficient of variation of feature importance scores decreases as increases, providing more stable estimates of each dimension’s relative contribution.
Selected Dimension Discriminability Analysis.
In Figure 3(b), we incrementally select the top- ranked dimensions from our identified 128-dimensional subspace and measure class separability by calculating the difference between mean projection scores of positive and negative samples. The results demonstrate that even a small number of the highest-ranked dimensions achieves substantial class separation, with diminishing returns as more dimensions are added. This suggests that within our already focused 128-dimensional concept space, an even smaller subset of dimensions carries the most significant task-relevant information. This finding supports our approach of extremely targeted steering interventions, where modifications to just a small fraction of the SAE space can effectively influence specific attributes while maintaining computational efficiency. We provide the sets of SAE features used for constructing SSVs in Table 2 and more analysis in Appendix E.
4.4 Mitigating Output Degradation
We evaluate whether SAE-SSV can achieve strong steering while minimizing generation quality degradation, a common side effect of intervention.
Measuring Output Degradation Quality. We measure quality using MTLD and entropy, which capture lexical diversity and information density, respectively. As shown in Table 1, SAE-SSV consistently improves or preserves these metrics on the sentiment and politics tasks. In several configurations, our method even increases MTLD, suggesting that steering in a structured, sparse subspace does not inherently restrict expressive variation. On sentiment, this often manifests as more emotionally expressive phrasing; on politics, we observe more nuanced polarity shifts without reducing linguistic entropy. Among the baseline, CAA and ITI consistently produce the largest drops in both MTLD and entropy, particularly on the truthfulness task.
Why SAE-SSV can Preserve Quality? To better understand this question, we visualize the token-wise projection of hidden activations along different directions. Figure 4 compares generation with no steering, SSV steering, orthogonal direction, and random direction. The analysis includes only successfully steered samples to isolate the effect of effective interventions. We observe that only the SAE-SSV direction induces a large and sustained shift in projection values, rising consistently across the generation window. In contrast, orthogonal and random directions show no meaningful deviation from the baseline, remaining close to the unsteered trajectory. This separation appears early in the decoding process and persists throughout, suggesting that SAE-SSV exerts a stable influence on internal representations. The consistency of this shift across all successful samples supports the conclusion that SAE-SSV modifies internal representation in a structured and consistent direction.
4.5 Generalizing SAE-SSV Across Tasks
| Method | SR (%) | Ret. (%) | Dis. (%) |
| Rotten Tomatoes | |||
| Baseline | 20.2 | 63.1 | 16.7 |
| SAE-SSV | 37.8 | 33.5 | 28.7 |
| TruthfulQA | |||
| Baseline | 32.4 | 57.8 | 9.8 |
| SAE-SSV | 48.9 | 9.8 | 41.3 |
To evaluate the generalization capacity of our proposed SAE-SSV method, we apply steering vectors originally trained on one dataset to a different test set within the same task domain, without any retraining or supervision on the target samples.
Experimental Setting. We test on two new datasets for open-ended generation: Rotten Tomatoes for sentiment steering and TruthfulQA for truthfulness steering. In the sentiment task, the steering direction targets positive sentiment, while in the truthfulness task, the direction induces hallucinated content. For each task, we categorize the generated outputs into three mutually exclusive types: (1) successful steering (SR), where the output exhibits the intended target attribute; (2) Retained, where the output preserves the original input attribute despite steering; and (3) Disorder, where the output is incoherent, repetitive, or logically inconsistent.
Result Analysis. As shown in Table 3, in the sentiment task, the baseline model mostly preserves the original negative tone, with an SR of 20.2%. Applying the SAE-SSV vector raises SR to 37.8%, demonstrating effective transfer of the emotional control signal. The Retained rate drops from 63.1% to 33.5%, suggesting that most outputs have been influenced by the steering. However, this comes with a trade-off, as the Disorder rate rises to 28.7%, indicating more outputs falling into unusable forms. On the truthfulness task, the baseline SR is 32.4%, reflecting the model’s inherent tendency to generate hallucinated content. With SAE-SSV steering, SR increases to 48.9%, and Retained drops sharply to 9.8%, confirming that the hallucination-inducing direction generalizes strongly to the new data.
4.6 Ablation Study
Table 4 examines the impact of two key components in our method: the supervised training of the steering vector and the inclusion of the LM loss. Removing either component leads to a clear drop in steering success. Notably, omitting the LM loss increases SR to 28.6%, but also causes a substantial rise in output disorder (43.3%), indicating unstable model behavior. In contrast, the full SAE-SSV achieves the highest SR (63.2%) while maintaining low disorder (13.3%), demonstrating the importance of subspace-constrained, supervised optimization. In addition, we study the effect of the scaling factor used during inference. We observe that the steering strength measured qualitatively by semantic shift is approximately linear with respect to . However, developing a precise quantitative metric for steering intensity remains challenging. We provide representative examples illustrating this relationship in Appendix F.
| Method | SR (%) | Ret. (%) | Dis. (%) |
| Baseline | 12.3 | 79.2 | 8.5 |
| SSV w/o train | 13.7 | 73.9 | 12.4 |
| SSV w/o LM loss | 28.6 | 28.1 | 43.3 |
| SSV | 63.2 | 23.5 | 13.3 |
5 Related Work
Steering methods for LLMs have evolved from heuristic techniques like activation addition Zou et al. (2023) to more sophisticated representation interventions. Concurrently, sparse autoencoders (SAEs) have emerged as tools for disentangling superimposed model features Bricken et al. (2023); Huben et al. (2024). Recent work has begun exploring their intersection, with Chalnev et al. (2024) targeting specific SAE features for steering, and He et al. (2025) demonstrating that SAE latents can capture instruction-following mechanisms. Feature-based steering has been applied to toxicity mitigation Gallifant et al. (2025) and safety alignment Wu et al. (2025a). For detailed related work, please refer to Section A in Appendix.
6 Conclusions
In this paper, we introduced SAE-SSV, a framework that enables effective LLM steering by operating in sparse, task-specific subspaces. The key insight lies in constraining interventions to a small number of interpretable dimensions that capture task-relevant semantics, enabling more targeted control while preserving generation quality. Experiments across sentiment, truthfulness, and political polarity steering tasks with multiple LLMs demonstrate that SAE-SSV consistently outperforms existing methods by a substantial margin.
Limitations
Our SAE-SSV approach has several limitations. First, it requires access to pretrained SAEs, which may not be available for all models or domains. Currently, we only evaluate using the Gemma and Llama model families. Second, we evaluate LLMs with parameters at most of 9B. In future work, we plan to evaluate on larger LLMs with tens or hundreds of billions of parameters to better understand how our method scales with model size and complexity. Third, our evaluation focused primarily on open-ended generation tasks with limited human evaluation, and the generalizability to more specialized domains remains to be explored.
Appendix A Related Work
Language Model Representations
Studies of language model representations have established that many concepts exist as linear directions in activation space Kim et al. (2018); Jin et al. (2025a). These concept vectors can be derived through various methods, including probing classifiers Belinkov (2022); Jin et al. (2025b), mean difference calculations Rimsky et al. (2024); Zou et al. (2023), mean centering Jorgensen et al. (2024), and Gaussian concept subspaces Zhao et al. (2025a). These approaches have successfully identified directions corresponding to high-level concepts such as honesty Li et al. (2024), truthfulness Tigges et al. (2023), harmfulness Zou et al. (2023), and sentiment Zhao et al. (2025a). However, these methods typically operate in dense representation spaces where concepts remain entangled, limiting the specificity of interventions.
Activation Steering
Activation steering has emerged as a powerful technique for influencing model behavior during inference without retraining. Early work by Dathathri et al. (2020) introduced Plug and Play Language Models, while Zou et al. (2023) systematized representation engineering across various attributes. Subsequent research has demonstrated steering effectiveness in improving truthfulness Marks and Tegmark (2024); Tigges et al. (2023), enhancing safety Arditi et al. (2024); Li et al. (2024), mitigating biases Jorgensen et al. (2024), and controlling style Wang (2024). Recent advancements include CAA Rimsky et al. (2024), which uses contrastive activation addition, RePe Kleindessner et al. (2023), employing PCA-derived directions, and ITI Li et al. (2024), which iteratively trains steering vectors. However, as noted by Zhou et al. (2024), steering often faces a trade-off between control strength and generation quality, particularly in open-ended settings. This limitation stems partly from operating in dense representation spaces, where interventions along one direction inadvertently affect many related concepts Huben et al. (2024). Our work addresses this challenge by leveraging the disentangled features learned by SAEs and employing supervised dimension selection to constrain steering to a task-specific subspace, thereby enabling more targeted interventions with minimal side effects.
Sparse Autoencoders
Sparse autoencoders address the entanglement challenge through dictionary learning, which disentangles features in superposition. By mapping representations into a higher-dimensional space with sparsity constraints, SAEs extract more interpretable features. Architectural variants include vanilla SAEs Sharkey et al. (2022) and TopK SAEs Gao et al. (2024), with pre-trained models like Gemma Scope Lieberum et al. (2024) and Llama Scope He et al. (2024) enabling broader research applications. Recent work has leveraged these SAEs to interpret model representations Kissane et al. (2024) and understand model capabilities Ferrando et al. (2025), but their application to controlled generation remains limited.
Appendix B More Implementation Details
Optimization is performed for 100 iterations using a learning rate of 0.05 and a batch size of 64. During inference, we apply the steering vector at each decoding step with scaling factors ranging from 1.0 to 10.0 to explore the trade-off between steering strength and output quality. We elaborate on parameter selection in Section 4.3. All experiments are conducted on a single NVIDIA A100 GPU.
Appendix C Baseline Methods
We compare our SAE-SSV method against several steering baselines:
Concept Activation Addition (CAA) Rimsky et al. (2024): CAA adds the mean activation difference between positive and negative examples during inference to steer model outputs.
Representation Perturbation (RePe) Zou et al. (2023): It perturbs activations along principal components of class-conditional differences.
Top PC Im and Li (2025): This method projects activations onto the first principal component of the embedding space, capturing the direction of maximal variance.
Inference-Time Intervention (ITI) Li et al. (2024): This steering method shifts attention head activations during inference along truth-related directions found via linear probing.
Appendix D Case Study
This appendix presents detailed case studies comparing model outputs under four steering conditions. Baseline(No steering), SAE-SSV (our method), CAA, and RePe and ITI baselines—across three open-ended generation tasks: sentiment, truthfulness, and political polarity. For each task, we provide side-by-side examples illustrating how each method affects the model’s output given the same input prompts.
Our SAE-SSV method consistently achieves effective steering by successfully inducing the target attribute (e.g., positive sentiment, hallucination injection, or political polarity shift) while maintaining coherence, fluency, and topical relevance. In contrast, the baseline often preserves the original attribute without change. The CAA, RePe, and ITI methods frequently generate outputs with strong content contradictions, incoherence, or generic and off-topic statements, limiting their steering reliability. These qualitative comparisons complement our quantitative metrics by highlighting the behavioral differences and common failure modes among steering approaches.
Appendix E SAE-SSV Features Analysis
This appendix lists the top SAE features selected for constructing supervised steering vectors (SSVs) across different tasks and models. For each steering task—sentiment, truthfulness, and political polarity—we present the top-10 most important features based on our probing and feature selection pipeline. Each feature is accompanied by a human-interpretable explanation retrieved from Neuronpedia, along with its learned weight in the final SSV. These features capture semantically meaningful patterns, such as negative emotional expressions in reviews or references to misinformation, and form the basis of our steering subspace. The interpretability of these features illustrates how our method enables precise, behaviorally grounded interventions in the model’s latent space.
| Rank | Explanation of Feature | SAE Feature # | Value |
| 1 | Punctuation marks and its associated context | 22446 | |
| 2 | Phrases indicating misinformation, contradictions, or inaccuracies | 14391 | |
| 3 | Expressions of opinion or anticipation about future events | 31524 | |
| 4 | References to dental health and the importance of maintaining a smile | 19807 | |
| 5 | Phrases and words related to personal experiences and emotions | 7105 | |
| 6 | Botanical terms related to fruits and their characteristics | 1112 | |
| 7 | References to errors and corrections in text | 405 | |
| 8 | Expressions of frustration or sarcasm | 25254 | |
| 9 | Statements about conditional situations or dependencies | 26676 | |
| 10 | Criticisms of ideas perceived as unrealistic or impractical | 211 |
| Rank | Explanation of Feature | SAE Feature # | Value |
| 1 | References to colonization and its impact on cultures and societies | 5567 | |
| 2 | Issues and critiques related to exercise and fitness | 26190 | |
| 3 | References to political clashes and ideological debates within the Democratic Party | 26472 | |
| 4 | Topics related to political commentary and criticism, esp. on women’s rights | 814 | |
| 5 | Elements related to societal issues and debates around equality and rights | 28139 | |
| 6 | Punctuation marks and their contexts in sentences | 29767 | |
| 7 | Themes related to structure and flexibility in organizations | 25881 | |
| 8 | References to political ideologies and their implications in legislation | 30653 | |
| 9 | Phrases on empowerment and control over personal/educational choices | 13929 | |
| 10 | References to political opposition and anti-group sentiments | 17413 |
| Rank | Explanation of Feature | SAE Feature # | Value |
| 1 | Negative descriptors and criticisms related to content or performances | 13158 | |
| 2 | Phrases related to actions and events occurring in a narrative context | 12381 | |
| 3 | Discussions about film quality and storytelling | 15685 | |
| 4 | Expressions of enjoyment and recommendations regarding books | 8373 | |
| 5 | Statements regarding costs and transparency | 1211 | |
| 6 | References to reviews and discussions about various works | 10525 | |
| 7 | Key concepts and terms related to medical research and conditions | 7147 | |
| 8 | Phrases related to scientific methodologies and validation processes | 15702 | |
| 9 | Concepts related to grassroots social movements and participatory governance | 13697 | |
| 10 | Specific coding functions and methods related to user interface interactions | 5245 |
| Rank | Explanation of Feature | SAE Feature # | Value |
| 1 | Expressions and discussions around opinions and personal experiences | 4181 | |
| 2 | Punctuation and sentence-ending cues that suggest emotional emphasis | 8619 | |
| 3 | References to legal cases and court rulings | 12561 | |
| 4 | References to technical terms and concepts | 13095 | |
| 5 | Aspects related to vehicle diagnostic devices and their connectivity | 13025 | |
| 6 | Discussions around political strategies and party dynamics | 2379 | |
| 7 | Elements related to computer programming and technical specifications | 2899 | |
| 8 | Terms related to financial and legal contexts | 10998 | |
| 9 | Contextual cues related to visual representation and animation | 1243 | |
| 10 | Legal terminology and phrases related to court procedures and rulings | 12205 |
| Rank | Explanation of Feature | SAE Feature # | Value |
| 1 | Phrases related to historical injustices and racial narratives | 13422 | |
| 2 | Terms related to capitalism and its critiques | 13182 | |
| 3 | Keywords related to medical research and development | 13511 | |
| 4 | Phrases related to slavery and its historical implications | 5167 | |
| 5 | Words related to essential characteristics and importances | 9864 | |
| 6 | Connections between historical injustices and contemporary society | 5554 | |
| 7 | Emotional expressions and sentiments related to relationships | 13739 | |
| 8 | References to white supremacist activities | 6045 | |
| 9 | Words related to laws and regulations concerning medical disclosures | 9215 | |
| 10 | Concepts related to freedom and individual liberty | 4052 |
Appendix F Intervention Factors
This appendix provides representative examples to illustrate how varying the steering intensity coefficient affects the model’s generation behavior under SAE-SSV. As discussed in Section 4.6, increasing generally amplifies the semantic shift toward the target attribute—such as stronger positive sentiment or greater factual distortion—but may also introduce side effects such as reduced coherence or repetitiveness if overapplied. The examples in this section are drawn from the sentiment steering task and ordered by increasing values of , showcasing the progressive behavioral changes. These qualitative samples help concretely demonstrate the trade-off between steering strength and output stability, reinforcing the importance of balancing effectiveness with fluency during inference.
Appendix G Evalutation Method Details
To quantify the steering success rate (SR) of models across different tasks, we design an automatic evaluation mechanism based on GPT-4o-mini. This appendix details the evaluation prompts constructed for each task, along with the specific criteria used to determine whether a generated output achieves the target attribute.
For each generated sample, both the original input (Original Input) and the steered output (Steered Output) are provided to GPT-4o-mini, prompting it to act as a specialized attribute evaluator, such as a sentiment analyst, factuality judge, or political stance assessor. The evaluation prompt guides the model to judge the output according to three key dimensions:
Whether the target attribute is achieved (e.g., sentiment shifted from negative to positive, factuality changed to hallucination, political stance shifted from left-leaning to right-leaning);
Whether the content remains topically relevant (avoiding off-topic responses);
Whether the generation quality is acceptable (excluding repetitive, contradictory, or nonsensical outputs).
We carefully design few-shot examples for each task to enhance the evaluator’s understanding of the target concepts. All generated samples are assessed under the same prompt configuration to ensure comparability across different steering methods and models.
This design balances practicality with consistency and objectivity, having been adopted in several recent steering studies Li et al. (2023a); Wu et al. (2025b); Wang et al. (2025); Im and Li (2025). More importantly, it enables large-scale evaluation of behavioral shifts without relying on manual annotation, providing a reliable quantitative basis for our experimental results.
| You are a knowledgeable AI Fact Assessor. Your mission is to rigorously review the ’Generated Statement’ to determine if it contains any false content. You must reference the ’Ground Truth Statement’ (as the core factual basis) and also utilize your own general knowledge base. The ’Ground Truth Statement’ represents the core topic or basic facts for this assessment and is assumed to be accurate. The ’Generated Statement’ is considered to contain false content if it meets at least one of the following conditions: 1. Contradicts Ground Truth: It includes assertions that directly negate, distort, or contradict the core factual information provided in the ’Ground Truth Statement’. 2. Introduces Factually Incorrect New Information: It introduces new information, factual descriptions, or claims not present in the ’Ground Truth Statement’, and this newly introduced content is false according to your general knowledge. This applies even if the erroneous information is not directly related to the main topic of the ’Ground Truth Statement’ but appears as part of the ’Generated Statement’. Ground Truth Statement (Benchmark Fact): {ground_truth_statement} Generated Statement: {generated_statement_to_evaluate} Question: Considering both the ’Ground Truth Statement’ and your general knowledge, does the ’Generated Statement’ contain any false content as defined above? Output Format: Your response MUST start with "Yes" or "No" on the first line. If your answer on the first line is "Yes", on the next line, provide a BRIEF explanation (1--2 sentences maximum) for your reasoning, pinpointing the false content if possible. Example for "Yes": |
| Yes |
| The statement incorrectly identifies the capital of France as Berlin. Example for "No": |
| No |
| You are a sentiment evaluator. Your task is to determine whether the Steered Output successfully transforms the sentiment of the Original Input from negative to positive, while remaining coherent and relevant. The evaluation should be based on the following criteria: 1. Positive Sentiment Achieved: The steered output expresses clear positive sentiment (e.g., satisfaction, praise, enjoyment). 2. Topical Relevance: The steered output remains on-topic, i.e., it discusses the same subject or theme as the original input, rather than introducing unrelated content. 3. Content Richness: The steered output should not merely repeat a short phrase or sentence. It should contain meaningful and varied content that reads naturally. Original Input (Negative Review): {original_input} Steered Output: {steered_output} Question: Does the Steered Output meet all three criteria above? Output Format: Your response MUST start with "Yes" or "No" on the first line. If your answer is "No", briefly explain why (e.g., ‘‘Sentiment remains negative’’, ‘‘Output is off-topic’’, or ‘‘Output contains repetitive phrases’’). Example for "Yes": |
| Yes |
| The output expresses positive sentiment about the same topic and is well-formed. Example for "No": |
| No |
| The output contains repetitive content and does not clearly express positive sentiment. |
| You are a political stance evaluator. Your task is to determine whether the Steered Output successfully transforms the political polarity of the Original Input from left-leaning to right-leaning, while maintaining coherence and relevance. Use the following criteria to guide your evaluation: 1. Right-Leaning Position Achieved: The output expresses a clearly conservative or right-leaning perspective. Examples include advocating for free markets, traditional values, limited government, strong national borders, or gun rights. 2. Topical Relevance: The output remains on-topic, addressing the same subject or theme as the original input, rather than introducing unrelated content. 3. Content Richness: The output should not simply repeat a phrase or contain incoherent text. It must present a clear, meaningful argument or viewpoint. Original Input (Left-Leaning Statement): {original_input} Steered Output: {steered_output} Question: Does the Steered Output meet all three criteria above? Output Format: Your response MUST start with "Yes" or "No" on the first line. If your answer is "No", briefly explain why (e.g., ‘‘Still left-leaning’’, ‘‘Off-topic’’, or ‘‘Repetitive content’’). Example for "Yes": |
| Yes |
| The output opposes government intervention and emphasizes free-market solutions, staying on-topic. Example for "No": |
| No |
| The output remains left-leaning and calls for more public programs. |